can you help me with this problem? PLEASE I REALLY NEED HELP WITH IT!!!!!!!
(7-i)+(9+11i)

thank you
sandy

Answer by venugopalramana(3286)

SEE THE FOLLOWING AND YOU CAN EASILY ANSWER YOUR PROBLEM..THE ANSWER IS 16+10i....

can u help me with this specific problem i dont understand it its 4(7-i)-5(2-6i) i would greatly appreciate it thanks
1 solutions
Answer 10001 by venugopalramana(381)on 2005-11-28 23:25:43 
4(7-i)-5(2-6i)
HOPE YOU KNOW COMPLEX NUMBERS...FOR ELEMENTARY UNDERSTANDING IT SUFFICES TO KNOW THAT WE DESIGNATE SQUARE ROOT OF -1 AS i WHICH IS AN IMAGINARY NUMBER..WHERE AS 1,2,3 ETC..ARE ALL REAL NUMBERS..WE CANT MIX THEM LIKE WE CANT ADD 2 APPLES TO 3 GRAPES AND SAY THAT THERE ARE 5 APPLES OR 5 GRAPES..WE CAN ONLY SAY THAT THERE ARE 2 APPLES AND 3 GRAPES..SO WE DEAL WITH REAL NUMBERS AND IMAGINARY NUMBERS SEPERATELY AND CALL THEM TOGETHER AS COMPLEX NUMBERS
4(7-i)-5(2-6i)=4*7-4i-(5*2-5*6i)=28-4i-10+30i=(28-10)+i(30-4)=18+26i